# Controllable Generation with Text-to-Image Diffusion Models: A Survey

### v1

@@ 2.3 Taxonomy
{forest}
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=left, font=, rectangle, draw=hidden-draw, rounded corners, align=left, minimum width=2em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, line width=0.8pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=6em,font=,, where level=2text width=4.3em,font=,, where level=3text width=4.0em,font=,, [ Controllable Generation with Text-to-Image Diffusion Models, ver [ Generation with
specific condition (§4) [ Personalization
(§ 4.1) [ (a) Subject-Driven: Textual Inversion[41] , Dreambooth[42] , Re-Imagen[43] , DreamArtist[44] ,
Custom Diffusion[45] , DVAR[46] , E4T[47] , ELITE[48] , UMM-Diffusion[49] , XTI[50] , SVDiff[51] ,
ANOVA[52] , SuTI[53] , Jia et al.[54] , InstantBooth[55] , COTI[56] , Gradient-Free TI[57] , Perfusion[58] ,
DisenBooth[59] , BLIP-Diffusion[60] , ProSpect[61] , Break-A-Scene[62] , COMCAT[63] , OFT[64] ,
PACGen[65] , Arar et al. [66] , Subject-Diffusion[67] , LyCORIS[68] , Kosmmos-G[69] , MCPL[70] ,
He et al.[71] , KC Loss[72] , MATTE[73] , Lego[74] , CatVersion[75] , CLiC[76] , VideoAssembler[77] ,
HiFi Tuner[78] , VideoBooth[79] , CAFE[80] , DETEX[81] , DreamTuner[82]
(b) Person-Driven: FastComposer[83] , Giambi et al.[84] , Face0[85] , DreamIdentity[86] ,
HyperDreamBooth[87] , PhotoVerse[27] , MagiCapture[88] , Face-diffuser[89] , $\mathscr{W}_{+}$ Adapter[90] ,
RetriBooru[91] , FaceStudio[92] , ViscoNet[93] , DemoCaricature[94] , PhotoMaker[95] , Stellar[96] ,
PortraitBooth[97]
(c) Style-Driven: StyleDrop[98] , StyleCrafter[99] , ArtAdapter[100] , StyleAligned[101] , SAG[102]
(d) Interaction-Driven: Reversion[103] , AnimateDiff[104] , MotionDirector[105] , LAMP[106] , SAVE[107] ,
Materzynska et al.[108] , DreaMoving[109] , MotionCrafter[110] , InteractDiffusion[111]
(e) Image-Driven: unCLIP[112] , Versatile Diffusion[113] , Prompt-Free Diffusion[114] , Uni-ControlNet[115] ,
IP-Adapter[116] , ViscoNet[93] , Context Diffusion[117] , FreeControl[118]
(f) Distribution-Driven: Cao et al. [119] , DreamDistribution[120] , leaf, text width=26.5em ] ] [ Spatial Control
(§ 4.2) [ eDiff-I[26] , LGP[121] , SpaText[122] , GLIGEN[123] , Universal Guidance[124] , LayoutDiffuse[125] ,
MCM[126] , FreeDoM[127] , FLIS[128] , LayoutDiffusion[129] , HumanSD[130] , LCDG[131] ,
Giambi et al.[84] , GeoDiffusion[132] , Attention-Refocusing[133] , ZestGuide[134] , CAC[135] , SSMG[136] ,
DenseDiffusion[137] , JointNet[138] , HyperHuman[139] , Region&Boundary[140] , EOCNet[141] ,
MATTE[73] , LoCo[142] , AnyLens[143] , LRDiff[144] , Loose Control[145] , DemoCaricature[94] ,
InteractDiffusion[111] , FreeControl[118] , Local Control[146] , SCEdit[147] , Ren et al. [148] , leaf, text width=26.5em ] ] [ Advanced
Text-Conditioned
(§ 4.3) [ Structure Diffusion[149] , Attend-and-Excite[150] , GlueGen[151] , Rich-text-to-image[152] , SynGen[153] ,
Tailored Visions[154] , ParaDiffusion[155] , PEA-Diffusion[156] , leaf, text width=26.5em ] ] [ In-Context (§ 4.4) [ Prompt Diffusion[157] , iPromptDiff[158] , leaf, text width=10.5em ] ] [ Brain-Guided (§ 4.5), text width=5.4em [ Mind-Vis[159] , Takagi et al. [160] , Brain-Diffuser[161] , MindDiffuser[162] , Ni et al. [163] ,
DreamDiffusion[164] , BrainVis[165] , leaf, text width=22.5em ] ] [ Sound-Guided (§ 4.6), text width=5.5em [ GlueGen[151] , CAVT-MAE[166] , leaf, text width=8.5em ] ] [ Text Rendering (§ 4.7), text width=5.5em [ Liu et al. [167] , GlyphDraw[168] , TextDiffuser[169] , GlyphControl[170] , AnyText[171] ,
TextDiffuser-2[172] , UDiffText[173] , Diff-Text[174] , leaf, text width=24.8em ] ] ] [ Generation with
[…]
Generation (§ 6) [ Conditional Score Prediction (§ 6.1), text width=8.5em [ DiffBlender[186] , Emu2[187] , leaf, text width=8.5em ] ] [ Condition-Guided Score Estimation (§ 6.2), text width=10.5em [ Universal Guidance[124] , FreeDoM[127] , SAG[102] , leaf, text width=14.5em ] ] ] ]
Figure 2: Taxonomy of Controllable Generation. From the condition perspective, we categorize controllable generation approaches into three sub-tasks, including generation with specific conditions, generation with multiple conditions, and universal controllable generation.
The task of conditional generation utilizing text-to-diffusion models represents a multifaceted and intricate domain. From the condition perspective, we divide this task into three sub-tasks (refer to Figure 2). Most works study how to generate images under specific conditions, e.g. image-guided generation, and sketch-to-image generation. To reveal the mechanical theory and features of these approaches, we further categorize them according to their condition types. The primary challenge in this task lies in how to enable pretrained text-to-image (T2I) diffusion models to learn to model new types of conditions and to generate in conjunction with textual conditions while ensuring the images produced are of high quality. Additionally, some methods investigate how to generate images using multiple conditions, such as given a character’s identity and pose. The main challenge in these tasks is the integration of multiple conditions, necessitating the capability to express several conditions simultaneously in the generated results. Furthermore, some works attempt to develop a condition-agnostic generation approach that can utilize these conditions to produce results.
@@
3 How to Control Text-to-Image Diffusion Models with Novel Conditions
In this section, we present the controlling mechanism of diffusion models from a score-based perspective [19]. Following [188], we can set the approximate denoising transition mean $\mu_{\theta}(x_{t},t)$ in Equation.3 as:
@@ 3 How to Control Text-to-Image Diffusion Models with Novel Conditions
where $s_{\theta}(x_{t},t)$ is a neural network that learns to predict the score function $\nabla_{x_{t}}\log p_{t}(x)$. In DDPM, we have:
@@
where $\epsilon\sim\mathcal{N}(0,\mathbf{I})$ is the gaussian noise used in forward process, $\alpha_{t}:=1-\beta_{t}$, and $\bar{\alpha}_{t}:=\prod_{s=0}^{t}\alpha_{s}$. Then, Equation.4 can be written as:
@@
In conditional generation ($c$ denotes condition), the score function is extended with a posterior probability term $\nabla_{x_{t}}\log p_{t}(c|x)$ and becomes $\nabla_{x_{t}}\log\left(p_{t}(x)p_{t}^{w}(x|c)\right)$ ($w$ represents a hyper-parameter to control condition intensity), following [18, 20]. To employ a neural network for conditional generation, classifier-free guidance (CFG)[20] transforms it to:
$\nabla_{x_{t}}\log$ $\left(p_{t}(x)p_{t}^{w}(x|c)\right)$
$=\nabla_{x_{t}}\log p_{t}(x)+w\nabla_{x_{t}}\log p_{t}(c|x)$
$=\nabla_{x_{t}}\log p_{t}(x)+w\nabla_{x_{t}}\log\frac{p_{t}(x|c)}{p_{t}(x)}$
$=(1-w)\nabla_{x_{t}}\log p_{t}(x)+w\nabla_{x_{t}}\log p_{t}(x|c)$ (7)
where $\nabla_{x_{t}}\log p_{t}(x)$ and $\nabla_{x_{t}}\log p_{t}(x|c)$ can be predicted by training a model $\epsilon_{\theta}(x_{t},\cdot,t)$, which predict the former via $\epsilon_{\theta}(x_{t},\phi,t)$ and the latter via $\epsilon_{\theta}(x_{t},c,t)$. Existing T2I diffusion models train $\epsilon_{\theta}(x_{t},\cdot,t)$ by randomly dropping the text prompt, and the denoising process with CFG is as follows:
@@
and $\hat{\epsilon}(x_{t},c_{text},t)$ is used in Equation.4 for conditional synthesis.
Hence, the key to controlling text-to-image models with novel conditions $c_{novel}$ is to model score $\nabla_{x_{t}}\log p_{t}(x|c_{text},c_{novel})$. Following [18, 32], there are two types of mechanisms, i.e., conditional score prediction and conditioned-guided score estimation, which we illustrate below.
@@ 3.1 Conditional Score Prediction
(a) Model-based Conditional Score Prediction.
(b) Tuning-based Conditional Score Prediction.
(c) Training-free Conditional Score Prediction.
Figure 3: Illustrations of conditional score prediction mechanisms. .
While T2I diffusion models leverage $\epsilon_{\theta}(x_{t},c_{text},t)$ to predict $\nabla_{x_{t}}\log p_{t}(x|c_{text})$, a fundamental and powerful way for steering diffusion models is through conditional score prediction in the sampling process, where these methods introduce $c_{novel}$ into $\epsilon_{\theta}(x_{t},c_{text},t)$, constructing a $\tilde{\epsilon}(x_{t},c_{text},c_{novel},t)$ to straightforwardly predict $\nabla_{x_{t}}\log p_{t}(x|c_{text},c_{novel})$. Then, the denoising process with CFG of conditional score prediction methods is as follows:
$\hat{\epsilon}(x_{t},c_{text},c_{cond},t)=(1-w)$ $\tilde{\epsilon}(x_{t},\phi,t)$
$+w$ $\tilde{\epsilon}(x_{t},c_{text},c_{cond},t)$ (9)
• Model-based Conditional Score Prediction. Some approaches employs an additional encoder $E$ to encode novel conditions and input the encoded features into $\epsilon_{\theta}$, where the conditional score prediction process is as follows:
$\tilde{\epsilon}(x_{t},c_{text},c_{novel},t)$ $=\epsilon_{\theta^{*}}(x_{t},c_{text},E(c_{novel}),t)$ (10)
where $E$ and $\theta^{*}$ are trainable. The schematic illustration is shown in Figure 3a.
• Tuning-based Conditional Score Prediction. Tuning-based methods typically focus on adapting to a specific condition, often in scenarios with limited data, such as single or few-shot examples. These methods achieve conditional prediction by transforming either the text condition $c_{text}$ or the model parameters $\theta$ into a form specific to the given condition, as shown in Figure 3b. This can be represented as:
@@
• Training-free Conditional Score Prediction. While the above techniques require a training process, some methods are designed in a training-free manner (refer to Figure 3c). They introduce conditions to control the generation directly through the intrinsic ability of the structure of UNet, such as modulating the cross-attention map to control the layout [135, 142] or introducing features of the reference image in self-attention to control the style [101].
3.2 Condition-Guided Score Estimation
Figure 4: Illustration of condition-guided score estimation.
Unlike conditional score prediction approaches predicting $\nabla_{x_{t}}\log p_{t}(x|c_{text},c_{novel})$, condition-guided estimation approaches are designed to gain $\nabla_{x_{t}}\log p_{t}(c_{novel}|x_{t})$ without need of CFG, which generally train an additional model with parameters $\varphi$ to predict the condition from latent or internal features, denoting as $p_{\varphi}(c_{novel}|x_{t})$. It can be utilized to attain $\nabla_{x_{t}}\log p_{t}(c_{novel}|x)$ via backpropagation, as illustrated in Figure 4. And the denoising process now reads:
$+\gamma\nabla_{x_{t}}\log p_{\varphi}(c_{novel}|x_{t})$ (12)
4 Controllable Text-to-Image Generation with Specific Conditions
Figure 5: Illustration of controllable text-to-image generation with specific conditions. The condition is marked in blue background. Examples are sourced from [48, 27, 100, 189, 112, 119, 190, 157, 155, 162, 151, 174].
Building upon the foundation of text-to-image diffusion models, introducing novel conditions to steer the generative process represents a complex and multifaceted task. In the following chapters, we review the existing methods of conditional generation according to the condition perspective, providing a comprehensive critique of their methodologies.
4.1 Personalization
Personalization task aims to capture and utilize concepts as generative conditions, which are not easily describable through text, from exemplar images for controllable generation. In this section, we provide an overview of these personalized conditions, categorizing them to offer a clearer understanding of their diverse applications and functionalities. We illustrate the results of personalization in Figure 5.
4.1.1 Subject-Driven Generation
In this section, we provide a detailed overview of subject-driven generation methods. The subject-driven generation task (also known as subject-centric personalization) is designed to produce visual content that retains the subjects of provided samples. In practice, many subject-driven generation methods are not confined to conditions specific to subject types; they often demonstrate a more universal capability. Thus, many of the approaches discussed in this chapter can be extended to a wider range of customized tasks. In summarizing these works, we adopt a broader perspective to showcase their general applicability as much as possible, aiming to facilitate a better understanding of their contributions and roles.
According to the controlling mechanism mentioned in Section 3, since all of these methods use conditional score prediction to introduce the conditions, we category them by their pipelines: tuning-based methods, which adapt model parameters or embeddings to cater to specific conditions; model-based methods, employing encoders to extract personalized conditions and feeding them into diffusion models; and training-free methods, which leverage external references to steer the generative process without the need of training.
• Tuning-based Personalized Score Prediction. A simple yet effective way to grasp concepts from provided samples involves selectively tuning a subset of parameters to reconstruct these concepts within text-to-image models, where the updated parameters are tailored to the desired concepts [44, 41, 42, 50, 61, 72, 77].

### v2

@@ 2.6 Taxonomy
Conditional generation with text-to-image diffusion models is inherently complex and can be broadly organized into three sub-tasks from the perspective of conditioning signals. The first and most studied direction augments pretrained diffusion models with novel conditions. We group these methods by their theoretical foundations, namely conditional score prediction and condition-guided score estimation. The main challenge is to flexibly inject new condition types alongside text prompts without sacrificing image quality. The second direction targets multi-condition control, such as combining a character’s identity with a specific pose. Methods are categorized by their technical strategies, including joint training, continual learning, weight fusion, attention-based integration, and guidance composition. Here, the core difficulty is to fuse multiple signals so that all conditions are faithfully expressed. The third direction pursues condition-agnostic generation, aiming to build unified frameworks that can robustly leverage diverse condition types across a wide range of inputs.
@@
3 Control Text-to-Image Diffusion Models with Novel Conditions
Following [168], we can set the approximate denoising transition mean $\mu_{\theta}(x_{t},t)$ in Eq. 3 as:
@@ 3 Control Text-to-Image Diffusion Models with Novel Conditions
where $s_{\theta}(x_{t},t)$ is a neural network that learns to predict the score function $\nabla_{x_{t}}\log p_{t}(x)$. Hence, we have:
@@
where $\epsilon\sim\mathcal{N}(0,\mathbf{I})$ is the Gaussian noise used in forward process, $\alpha_{t}:=1-\beta_{t}$, and $\bar{\alpha}_{t}:=\prod_{s=0}^{t}\alpha_{s}$. Then, Eq. 9 can be written as:
@@
In conditional generation ($c$ denotes condition), the score function is extended with a posterior probability term $\nabla_{x_{t}}\log p_{t}(c|x_{t})$ and becomes $\nabla_{x_{t}}\log\left(p_{t}(x_{t})p_{t}^{w}(x_{t}|c)\right)$ ($w$ represents a hyper-parameter to control condition intensity), following [57, 93]. To employ a neural network for conditional generation, classifier-free guidance (CFG)[93] transforms it to:
$\nabla_{x_{t}}\log$ $\left(p_{t}(x_{t})p_{t}^{w}(c|x_{t})\right)$
$=\nabla_{x_{t}}\log p_{t}(x_{t})+w\nabla_{x_{t}}\log p_{t}(c|x_{t})$
$=\nabla_{x_{t}}\log p_{t}(x_{t})+w\nabla_{x_{t}}\log\frac{p_{t}(x_{t}|c)}{p_{t}(x_{t})}$
$=(1-w)\nabla_{x_{t}}\log p_{t}(x_{t})+w\nabla_{x_{t}}\log p_{t}(x_{t}|c)$ (12)
where $\nabla_{x_{t}}\log p_{t}(x_{t})$ and $\nabla_{x_{t}}\log p_{t}(x_{t}|c)$ can be predicted by training a model $\epsilon_{\theta}(x_{t},\cdot,t)$, which predict the former via $\epsilon_{\theta}(x_{t},\phi,t)$ and the latter via $\epsilon_{\theta}(x_{t},c,t)$.
Existing T2I diffusion models train $\epsilon_{\theta}(x_{t},\cdot,t)$ by randomly dropping the text prompt, and the denoising process with CFG is as follows:
@@
and $\hat{\epsilon}(x_{t},c_{text},t)$ is used in Eq. 9 for conditional synthesis.
Hence, the key to controlling text-to-image models with novel conditions $c_{novel}$ is to model score $\nabla_{x_{t}}\log p_{t}(x_{t}|c_{text},c_{novel})$. Following [57, 318], there are two types of mechanisms, i.e., conditional score prediction and conditioned-guided score estimation, which we illustrate below.
• Conditional Score Prediction (Sec. 3.1). While T2I diffusion models leverage $\epsilon_{\theta}(x_{t},c_{text},t)$ to predict $\nabla_{x_{t}}\log p_{t}(x_{t}|c_{text})$, a fundamental and powerful way for steering diffusion models is through conditional score prediction in the sampling process, where these methods introduce $c_{novel}$ into $\epsilon_{\theta}(x_{t},c_{text},t)$, constructing a $\tilde{\epsilon}(x_{t},c_{text},c_{novel},t)$ to straightforwardly predict $\nabla_{x_{t}}\log p_{t}(x_{t}|c_{text},c_{novel})$. The CFG-based denoising update then reads:
$\hat{\epsilon}(x_{t},c_{text},c_{novel},t)=(1-w)$ $\tilde{\epsilon}(x_{t},\phi,t)$
$+w$ $\tilde{\epsilon}(x_{t},c_{text},c_{novel},t)$ (14)
• Condition-guided Score Estimation (Sec. 3.2). Unlike conditional score prediction straightforwardly predicting $\nabla_{x_{t}}\log p_{t}(x_{t}|c_{text},c_{novel})$, condition-guided estimation approaches estimate $\log p_{t}(c_{novel}|x_{t})$ with a likelihood/critic and obtain $\nabla_{x_{t}}\log p_{t}(c_{novel}|x_{t})$ by backpropagation, which is then injected into the sampler. And the denoising process now reads:
$+\gamma\nabla_{x_{t}}\log p_{t}(c_{novel}|x_{t})$ (15)
@@ 3.1 Conditional Score Prediction
Conditional score prediction approaches focus on empowering pre-trained denoising models supporting novel conditions to straightforwardly predict denoised latents. According to mechanisms, these methods can be categorized into tuning-based (Sec. 3.1.1), adapter-based (Sec. 3.1.2), and training-free (Sec. 3.1.3) manners.
Figure 3: Illustration of tuning-based conditional score prediction.
3.1.1 Tuning-based Conditional Score Prediction
Tuning-based methods typically focus on adapting to a specific condition, often in scenarios with limited data, such as single or few-shot examples. These methods achieve conditional prediction by transforming either the text condition $c_{text}$ or the model parameters $\theta$ into a form specific to the given condition, as shown in Fig. 3. This can be represented as:
@@ 3.1.1 Tuning-based Conditional Score Prediction
• Basic Tuning-based Methods. A straightforward yet effective strategy for learning controlling concepts from given samples is to fine-tune the diffusion models, thereby adapting text-to-image models to reflect the target conditions. The updated parameters are specialized to capture the desired conditions [59, 68, 221, 262, 328, 220, 333, 25].